OpenStudy (anonymous):

If a coin is flipped 10 times, what is the probability that it will show all heads or all tails? 1/2 1/512 1/1,022 1/1,024 1/2,048

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

1/2 on everyone chance. But there are 10 chances to get only heads or only tails. So multiply 1/2 by itself 10 times.

OpenStudy (anonymous):

or (1/2)^10

OpenStudy (misty1212):

the denominator should be \(2^{10}\) for each all heads \(\frac{1}{2^{10}}\) all tails \(\frac{1}{2^{10}}\) add and get \[\frac{2}{2^{10}}=\frac{1}{2^9}\]

OpenStudy (anonymous):

I got a crazy decimal.

OpenStudy (misty1212):

lol look at your answer choices see any decimals?

OpenStudy (anonymous):

because its a really small fraction

OpenStudy (misty1212):

compute \(2^9\) put that in the denominator

OpenStudy (anonymous):

Ohh so it's c?

OpenStudy (misty1212):

usually is C

OpenStudy (anonymous):

xD Thank you :)

OpenStudy (misty1212):

but not this time!!

OpenStudy (anonymous):

But i just checked it.

OpenStudy (misty1212):

\[2^{10}=1024\] but you want \[2^9\]

OpenStudy (misty1212):

all heads OR all tails two choices

OpenStudy (anonymous):

oh, so what do I do? 2^9?

OpenStudy (anonymous):

@misty1212 ?

OpenStudy (anonymous):

I think she is saying put 2^9 as the denominator.

OpenStudy (anonymous):

Okay, so when I did 2^9 in my calculator, i got 512, so would the answer be 1/512?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (anonymous):

No problem